OpenStudy (anonymous):

Determine the number of solutions for the quadratic function f(x) = x2 - 6x + 9. (Points : 2) No real solutions 1 solution (double root) 2 distinct real solutions

OpenStudy (mertsj):

Calculate the discriminant. Do you know what that is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1 solution discriminant = 0 x=3 is solution

OpenStudy (mertsj):

It is the part of the quadratic formula that is under the radical. b^2-4ac

OpenStudy (mertsj):

Way to go jadest

OpenStudy (anonymous):

thanks guys!

OpenStudy (mertsj):

yw

OpenStudy (across):

I believe this kind of questions tend to confuse students down the road; every polynomial of the form \(a_nx^n+a_{n-1}x^{n-1}+...+a_1x+a_0=0\) has \(n\) solutions in \(\mathbb{C}\). It seems teachers nowadays refer to "solutions" as those who can only be found in \(\mathbb{R}\). Someone correct me if I'm wrong.\[\]